OpenStudy (anonymous):

In a game where you can earn positive and negative points, you made these points on the first 4 plays: -5, -6, 8, 10. What is your current score?

OpenStudy (anonymous):

Do you want an explanation or just the answer? (:

OpenStudy (anonymous):

\[-5+ (-6)+8+10\]

OpenStudy (anonymous):

10 + 8 - 6 -5 = 7

OpenStudy (anonymous):

can you do both? i got the answer but i just wanted to see if i did it right

OpenStudy (anonymous):

Well, 10 + 8 is 18. Since there is a negative 5 and 6, you need to subtract 5 and 6 from 18. In the end you get positive 7 as your answer (:

OpenStudy (anonymous):

Because 18 - 6 = 12 and 12 - 5 = 7 (:

OpenStudy (anonymous):

Thanks so much! c:

OpenStudy (anonymous):

Of course (: